Evaluate the integral of √(3x + 1) dx from 1 to 5


1
Expert's answer
2021-03-24T14:21:03-0400

Solution.

"\int\limits_1^5\sqrt{3x+1}dx."

Make a substitution:

"t=3x+1," then "dt=3dx," hence "dx=\frac{1}{3}dt."

"\int\sqrt{3x+1}dx=\int\frac{1}{3}\sqrt{t}dt=\newline
=\frac{1}{3}\frac{t^{\frac{3}{2}}}{\frac{3}{2}}=\frac{2}{9}t^{\frac{3}{2}}=\newline
=\frac{2}{9}(3x+1)^{\frac{3}{2}}."

So,

"\int\limits_1^5\sqrt{3x+1}dx=(\frac{2}{9}(3x+1)^{\frac{3}{2}})|_1^5=\newline
=\frac{128}{9}-\frac{16}{9}=\frac{112}{9}=12\frac{4}{9}."

Answer: "\int\limits_1^5\sqrt{3x+1}dx=12\frac{4}{9}."